Sol Summary – January 27th

MDRS Sol Summary Report for Sol 12
Crew Journalist & Artist Niamh Shaw

Summary Title
The curse of the Gigabyte Hobgoblin

Mission Status:
We are all doing well, our experiments are now complete and we are
finishing off our remaining out-reach activities as part of our mission.

Sol Activity Summary:
Action! So many movies were made today it was hard to keep up with all the
activity. And the weather couldn’t have been better for shooting. But first
we took a call from The Sunday Times newspaper for an interview. We thought
it was going to be a brief 20 minutes group call and ended up being over 90
minutes in duration. What can I say? We wowed him with stories of our
mission. Here’s a brief sum-mary of the outreach and art activities
happening around here today: Niamh was shooting her art pro-ject in the
communal areas, Roy was filming his outreach movies in the Science Dome,
Idriss was using the EVA to make an outreach movie and Rick pulled himself
in many directions to assist every-one in their needs. Roy then made a
‘walk and talk’ outreach movie about MDRS and our life here. Michaela and
Rick went out on a second EVA of the day to collect one last set of
hypolith samples. Michaela then spent some time in the Science Dome
analyzing her samples and working on her scien-fitic projects. Lastly Roy
and Idriss finally built their wall with 3d printed bricks, and filmed
their suc-cessful proof of concept. We took over 600 photographs today and
made approx 10 movies. All content for us to bring back to our home
countries and share our adventure with everyone. On the home stretch now….

Look Ahead Plan:
We have to finish some minor outreach stuff in the morning, give the Hab a
good clean for Crew 174 and head out on one last EVA as a crew together.
Then its one last group picture and then we look forward to welcoming the
new crew to MDRS.

Anomalies in work:
The gigabyte goblin continues to strike, we will never know who or what was
eating all that data.

A perfectly wonderful day on Mars today. Warm, the sun shone and dry all

Crew Physical Status:
All crew are alll doing well. Still laughing.

A short EVA in the morning for Niamh and Idriss for outreach activities and
then in the afternoon Michaela and Rick went on EVA to collect hypolith

Reports to be filed:
sol 12 summary
Eva report
Eva request
commander’s report
Operations report